Use a calculator to help solve the problem. If the answer is not exact, give it to the nearest tenth.
The legs of a right triangle are equal, and the hypotenuse is 8.48528 units long. Find the length of each leg.

Answer by Alan3354(69443)

The legs of a right triangle are equal, and the hypotenuse is 8.48528 units long. Find the length of each leg.
------------
c^2 = a^2 + a^2 = 8.48528^2